OpenStudy (anonymous):

Please answer me question. 3m^2 - 3m / 4m^2 - 4m = 3/4, m does not equal to zero. Why is its restrictions m does not equal to zero?

OpenStudy (anonymous):

In addition is it 3m(m-1)/4m(m-1) = 3/4 m, m\[\neq\]0

OpenStudy (agreene):

the reason is: 3(0^2)-3(0)/4(0^2)-4(0) = 0/0 = undefined. and undefined =/= 3/4